you are asked to prepare a pH= 9.43 buffer starting from 60.0mL of 0.10M solution of ammonia, NH3, And 0.1M NH4Cl. (Kb for NH3= 1.8x10^-5)
How many mL of NH4Cl should be added to prepare buffer solution?
Answer #1

pH = 9.43

pOH = 14 - pH

pOH = 14 - 9.43

pOH = 4.57

Kb = 1.8 x 10-5

pKb = -log(Kb) = -log(1.8 x 10-5) = 4.74

moles NH3 = (molarity NH3) * (volume NH3)

moles NH3 = (0.10 M) * (60.0 mL)

moles NH3 = 6.0 mmol

According to Henderson-Hasselbalch equation

pOH = pKb + log([conjugate acid] / [weak base])

pOH = pKb + log(moles NH4+ / moles NH3)

4.57 = 4.74 + log(moles NH4+ / 6.0 mmol)

log(moles NH4+ / 6.0 mmol) = 4.57 - 4.74

log(moles NH4+ / 6.0 mmol) = -0.17

moles NH4+ / 6.0 mmol = 10-0.17

moles NH4+ / 6.0 mmol = 0.67

moles NH4+ = 6.0 mmol * 0.67

moles NH4+ = 4.0 mmol

volume NH4Cl = (moles NH4+) / (molarity NH4Cl)

volume NH4Cl = (4.0 mmol) / (0.1 M)

volume NH4Cl = 40.1 mL